OpenStudy (anonymous):

solve for g 5g+4(-5+3g)=1-g

OpenStudy (tamtoan):

what is 4(-5 + 3g) = ?

OpenStudy (anonymous):

1-g

OpenStudy (tamtoan):

how do you get 1 - g? ...remember 4 multiply that whole thing in the parenthesis

OpenStudy (anonymous):

5g+3g+g=1+20

OpenStudy (anonymous):

9g=21 g=7/3